# Towards Effective Multiple-in-One Image Restoration: A Sequential and Prompt Learning Strategy

### 07 Jan 2024 v1

@@ A Degradation Settings of MiO IR
where $\beta$ is the scattering coefficient of the atmosphere, and $d\left(y\right)$ is the distance between the object and the camera. To obtain haze and haze-free image pairs, we first estimate the depth map and then sample the value of $\beta$ and $A$ to generate haze images with different degrees. Following [23], we set $A$ from 0.8 to 1, $\beta$ from 0.5 to 2.5 for the training data and the test data in In-Dis, and set $A$ from 0.8 to 1, $\beta$ from 2.5 to 3 for the test data in Out-Dis.
@@
Single In-Dis SR Blur Noise JPEG Rain Haze Low-Light Avg. Ipv.
Table A.1: PSNR results on In-Dis test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, respectively. ‘-EP’ means explicit prompt learning. ‘-S+EP’ means using sequential learning and explicit prompt learning together. ‘-Single’ means that the model is trained on the corresponding single task.
B Single IR Performance of Backbones
We also present the performance of each backbone on different single IR tasks as a reference. For each backbone network, we train 7 single IR task models and test them on the corresponding single task. We train SRResNet with 500K iterations, and train SwinIR, Restormer and Uformer with 200K iterations. The results are shown in Tab. A.1, where ‘-Single’ means that the model is trained on the corresponding single task. The results of MiO IR models are also shown as a reference.
From Tab. A.1, we can see that the results of single task models are better than that of the MiO IR models. This makes sense because they are trained individually for each task, resulting in 7 models for 7 tasks, while there is only one shared model for the 7 tasks in MiO IR. It is worth noting that when our sequential and prompt learning is adopted, the gap between MiO IR and single task IR is greatly reduced compared with mixed learning baseline, even by half on several backbones. This again demonstrates the effectiveness of our strategies. With the future development of MiO IR network design and learning strategy, MiO IR models may tie, even surpass single task models.
C More Visualization Results
We have provided the visual results of competing methods on the In-Dis test sets in the main paper. In this section, we show the visual results on Out-Dis and Unknown test sets in Fig. C.1 and Fig. C.2, respectively. Though on the unseen data sets, all models show relatively lower performance, we can still see that models trained by our sequential and prompt learning strategies could achieve better visual results than that trained by mixed learning.
Figure C.1: Visual comparison of the results by different models on the 7 Out-Dis MiO test sets. (Zoom in and follow the arrows for the best view).
Figure C.2: Visual comparison of the results by different models on the 5 Unknown MiO test sets. (Zoom in and follow the arrows for the best view).

### 17 Mar 2024 v2

@@ B Degradation Settings of MiO IR
Figure B.1: Deraining model removes rain but retains JPEG artifacts. MiO model removes them simultaneously but obtains a lower PSNR because the GT images in the current deraining dataset contain JPEG artifacts. Therefore, simply adopting the datasets of single IR tasks is inappropriate for the investigation of MiO IR tasks. Please zoom in for better view.
@@
Single In-Dis SR Blur Noise JPEG Rain Haze Low-Light Avg. Ipv.
Table B.1: PSNR results on In-Dis test sets. ‘-M’ and ‘-S’ mean mixed and sequential learning, respectively. ‘-EP’ means explicit prompt learning. ‘-S+EP’ means using sequential learning and explicit prompt learning together. ‘-Single’ means that the model is trained on the corresponding single task.
@@
where $\beta$ is the scattering coefficient of the atmosphere, and $d\left(y\right)$ is the distance between the object and the camera. To obtain haze and haze-free image pairs, we first estimate the depth map (following [29]) and then sample the value of $\beta$ and $A$ to generate haze images with different degrees. Following [24], we set $A$ from 0.8 to 1, $\beta$ from 0.5 to 2.5 for the training data and the test data in In-Dis, and set $A$ from 0.8 to 1, $\beta$ from 2.5 to 3 for the test data in Out-Dis.
@@
C Single IR Performance of Backbones
We also present the performance of each backbone on different single IR tasks as a reference. For each backbone network, we train 7 single IR task models and test them on the corresponding single task. We train SRResNet with 500K iterations, and train SwinIR, Restormer and Uformer with 200K iterations. The results are shown in Tab. B.1, where ‘-Single’ means that the model is trained on the corresponding single task. The results of MiO IR models are also shown as a reference.
From Tab. B.1, we can see that the results of single task models are better than that of the MiO IR models. This makes sense because they are trained individually for each task, resulting in 7 models for 7 tasks, while there is only one shared model for the 7 tasks in MiO IR. It is worth noting that when our sequential and prompt learning is adopted, the gap between MiO IR and single task IR is greatly reduced compared with mixed learning baseline, even by half on several backbones. This again demonstrates the effectiveness of our strategies. With the future development of MiO IR network design and learning strategy, MiO IR models may tie, even surpass single task models.
D More Visualization Results
We have provided the visual results of competing methods on the In-Dis test sets in the main paper. In this section, we show the visual results on Out-Dis and Unknown test sets in Fig. D.2 and Fig. D.3, respectively. Though on the unseen data sets, all models show relatively lower performance, we can still see that models trained by our sequential and prompt learning strategies could achieve better visual results than that trained by mixed learning.
Figure D.2: Visual comparison of the results by different models on the 7 Out-Dis MiO test sets. (Zoom in and follow the arrows for the best view).
Figure D.3: Visual comparison of the results by different models on the 5 Unknown MiO test sets. (Zoom in and follow the arrows for the best view).